# SelaVPR++: Towards Seamless Adaptation of Foundation Models for Efficient Place Recognition

### v1

@@ V Discussion
Discussion on corner cases: Although SelaVPR++ is highly robust against most challenges in VPR, there are a few corner cases that it struggles to tackle effectively. Fig. 8 shows some examples of corner cases. For some extreme viewpoint changes or illumination changes (or even both simultaneously), other methods fail while ours shows good robustness and yields correct results. Besides, when the query image taken at natural scenes lacks discriminative landmarks, our method can still capture other useful details to identify places. However, our method fails in a few natural scenes. One solution is to increase the descriptor dimension to provide more detailed information, which needs to make a reasonable trade-off according to our needs. A more difficult case is that the vegetation changes drastically over time. This may require the model to have good reasoning ability to determine whether the change is reasonable in order to get the right result. In another failure case, the retrieved database image and the query are collected from two close locations and represent almost the same content, but the geographical distance between them exceeds the set threshold. This is considered a wrong match in VPR and is encountered by all methods. We believe that a direct mitigation way is to increase the geographical density of image collection when building the database.
Figure 8: Some examples of corner cases (success cases in the first three and failure cases in the last three). For the first example, the query is rotated 180. All other methods return the inverted but wrong results, while only our SelaVPR++ gets the correct image (with normal orientation). The second example shows a nighttime query with weak light and limited visibility. Our method is highly robust against light changes and gets the right database image (exhibiting obvious viewpoint changes with the query), but others fail. The third, fourth, and fifth query images are taken in nature, lacking discriminative landmarks and buildings. Our approach, which can capture other valuable details (e.g., vegetation), succeeds in the third case. However, this does not guarantee a correct match every time, especially when the plants shown in the query image have undergone significant growth or aging over time compared to the database image (i.e., the fourth example). In the last example, most methods return the same result that is geographically close to the query but exceeds the predefined threshold (i.e., failed). Although SALAD-CM introduces a novel CliqueMining strategy explicitly tailored to address this issue, it still yields the same wrong result as ours.
@@ Acknowledgments
This work was supported by the National Key R&D Program of China (2022YFB4701400/4701402), SSTIC Grant (KJZD20230923115106012, KJZD20230923114916032, GJHZ20240218113604008), Beijing Key Lab of Networked Multimedia, the Project of Pengcheng Laboratory (PCL2023A08, PCL2024Y02), and National Natural Science Foundation of China (62402252).

### v2

@@ V Discussion
Discussion on corner cases: Although SelaVPR++ is highly robust against most challenges in VPR, there are a few corner cases that it struggles to tackle effectively. Fig. 8 shows some examples of corner cases. For some extreme viewpoint changes or illumination changes (or even both simultaneously), other methods fail while ours shows good robustness and yields correct results. Besides, when the query image taken at natural scenes lacks discriminative landmarks, our method can still capture other useful details to identify places. However, our method fails in a few natural scenes. One promising solution is to increase the descriptor dimension to provide more detailed information, which requires making a reasonable trade-off according to our needs. In another failure case, the retrieved database image and the query are collected from two close locations and represent almost the same content, but the geographical distance between them exceeds the set threshold. This is considered a wrong match in VPR and is encountered by all methods. We believe that a direct mitigation way is to increase the geographical density of image collection when building the database.
@@ Acknowledgments
This work was supported by the National Key R&D Program of China (2022YFB4701400/4701402), SSTIC Grant (KJZD20230923115106012, KJZD20230923114916032, GJHZ20240218113604008), the Key Project of Pengcheng Laboratory, and National Natural Science Foundation of China (62402252).